Solve the partial differential equation (3D² +10DD'+ 3D ^1^2)z^ =e^x-y

1
Expert's answer
2022-01-06T11:28:45-0500

Since, the given equation is unclear, we assume it as:

"\quad Solve \left(D^{2}+4 D+3\right) y=e^{-3 x}"

Solution: Given equation is

"\left(D^{2}+4 D+3\right) y=e^{-3 x}"

A E is "\quad D^{2}+4 D+3=0"

"\begin{aligned}

&(D+1)(D+3)=0 \quad D=-1, \quad D=-3 \\

\therefore \quad & C F=c_{1} e^{-x}+c_{2} e^{-3 x} \\

& P I=\frac{1}{(D+1)(D+3)} e^{-3 x}

\end{aligned}"

put D=-3 in f(D) then

"\begin{aligned}

P I &=\frac{1}{0} e^{-3 x} \quad \therefore \text { method fails } \\

\therefore \quad P I &=\frac{e^{-3 x}}{(-3+1)(D-3+3)}\{1\}=\frac{e^{-3 x}}{(-2)} \frac{1}{D}\{1\}=\frac{e^{-3 x}}{-2} x \\

G . S . &=C F+P I

\end{aligned}"

"y=c_{1} e^{-x}+c_{2} e^{-3 x}-\frac{e^{-3 x}}{2} x"